OpenStudy (anonymous):

What are the diameter and radius of a circle with a circumference of 28π m? A. d = 28 m r = 14 m B. d = 14 m r = 28 m C. d = 14 m r = 7 m D. d = 7 m r = 14 m

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

\[\pi \div circumfronce\] That is how I remember you do it

OpenStudy (aaronandyson):

The circumference of a circle is C = 2*pi*r In this case, C = 28pi (given)

OpenStudy (aaronandyson):

plug that in to get C = 2*pi*r 28pi = 2*pi*r now solve for r and tell me what you get

OpenStudy (aaronandyson):

Here's how you solve for the radius r 28pi = 2*pi*r 28pi/pi = 2*pi*r/pi ... Divide both sides by pi. 28 = 2*r 28/2 = 2r/2 ... Divide both sides by 2. 14 = r r = 14

OpenStudy (anonymous):

Formula: \(C = 2 \pi d\) Plug in 28pi: \(28 \pi = \pi d\) Dividing pi to both sides gives us: \(28 = d\) So the diameter is 28, to find the radius we divide 28 / 2. And what's that? @torifame

OpenStudy (anonymous):

@torifamelol